(1-cosa)(1-cosb)(1-cosc)=sinasinbsinc find the value of (1=cosa)(1=cosb)(i+cosc)

Grade:11

7 Answers

Lab Bhattacharjee
8 years ago
\text{HINT: }\\ (1+\cos a)(1-\cos a)(1+\cos b)(1-\cos b)(1+\cos c)(1-\cos c)=\sin^2a\sin^2b\sin^2c
Lab Bhattacharjee
8 years ago
\text{HINT: }\\ (1+\cos a)(1-\cos a)(1+\cos b)(1-\cos b)(1+\cos c)(1-\cos c)=\sin^2a\sin^2b\sin^2c
Lab Bhattacharjee
8 years ago
\text{HINT: }\\ (1+\cos a)(1-\cos a)(1+\cos b)(1-\cos b)(1+\cos c)(1-\cos c)=\sin^2a\sin^2b\sin^2c
Lab Bhattacharjee
8 years ago
\text{HINT: }\\ (1+\cos a)(1-\cos a)(1+\cos b)(1-\cos b)(1+\cos c)(1-\cos c)=\sin^2a\sin^2b\sin^2c
Lab Bhattacharjee
8 years ago
\text{HINT: }\\ (1+\cos a)(1-\cos a)(1+\cos b)(1-\cos b)(1+\cos c)(1-\cos c)=\sin^2a\sin^2b\sin^2c
Lab Bhattacharjee
8 years ago
To the administrator, please delete the repeated answer(s).
 
Raghu Vamshi Hemadri
8 years ago
\text{HINT: }\\ (1+\cos a)(1-\cos a)(1+\cos b)(1-\cos b)(1+\cos c)(1-\cos c)=\sin^2a\sin^2b\sin^2c